OpenStudy (anonymous):

find the slope if it exists, of the line containing the pair of points. (-5,-4), and (-6,-10). The slope i s?

OpenStudy (anonymous):

m = y2 -y1/ x2 -x1 m = -10- 4/-6+5 m = -14/-1 m = 14 Slope is 14

OpenStudy (anonymous):

oh ok you set them up for x and y

OpenStudy (anonymous):

Mhm x1 and y1 and x2 and y2

OpenStudy (anonymous):

oh ok i was looking at it to set up similar to that but i did not know which number to put to subtract ....